### v1

# LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks

Abstract
Large language models (LLMs) have demonstrated impressive multilingual understanding and reasoning capabilities, driven by extensive pre-training multilingual corpora and fine-tuning instruction data. However, a performance gap persists between high-resource and low-resource language tasks due to language imbalance in the pre-training corpus, even using more low-resource data during fine-tuning. To alleviate this issue, we propose LinguaLIFT, a two-stage instruction tuning framework for advancing low-resource language tasks. An additional language alignment layer is first integrated into the LLM to adapt a pre-trained multilingual encoder, thereby enhancing multilingual alignment through code-switched fine-tuning. The second stage fine-tunes LLM with English-only instruction data while freezing the language alignment layer, allowing LLM to transfer task-specific capabilities from English to low-resource language tasks. Additionally, we introduce the Multilingual Math World Problem (MMWP) benchmark, which spans 21 low-resource, 17 medium-resource, and 10 high-resource languages, enabling comprehensive evaluation of multilingual reasoning. Experimental results show that LinguaLIFT outperforms several competitive baselines across MMWP and other widely used benchmarks.  [^1] Our code and data will be released once accepted.
LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks
Hongbin Zhang†‡,Kehai Chenthanks: Corresponding Author,Xuefeng Bai,Yang Xiang,Min Zhang Institute of Computing and Intelligence, Harbin Institute of Technology, Shenzhen, China Peng Cheng Laboratory, Shenzhen, China azure.starzhang@gmail.com,{chenkehai,baixuefeng,zhangmin2021}@hit.edu.cn, xiangy@pcl.ac.cn
1 Introduction
Large Language Models (LLMs) have demonstrated impressive understanding and reasoning capabilities, opening up a new paradigm for artificial intelligence (Wei et al., 2022; Zhao et al., 2023). These capabilities arise from self-supervised pre-training on large-scale multilingual data, followed by supervised fine-tuning on multilingual instruction data (Ouyang et al., 2022; Muennighoff et al., 2023; Chung et al., 2024). As a result, LLMs have achieved remarkable performance across various natural language processing tasks (Chen et al., 2024a, b; Zhang et al., 2024b, a), especially high-resource language mathematical reasoning tasks (Shi et al., 2023; She et al., 2024) and natural language inference tasks (Üstün et al., 2024; Lu et al., 2024).
Figure 1: Examples from the MGSM (Shi et al., 2023) dataset, where the mathematical problems share the same meaning across languages, but LLMs generate different answers. The red text marks the erroneous reasoning in the responses. Translations of the responses are provided in dashed boxes.
However, a substantial performance gap still exists between high-resource and low-resource language tasks (Zhu et al., 2024c), primarily due to the language imbalance within the large-scale multilingual corpora during pre-training, where high-resource languages dominate (Touvron et al., 2023). This gap remains unaddressed, even using more low-resource language data during the following fine-tuning Singh et al. (2024). As illustrated in Figure  1, while LLMs can generate the correct answer to the query in high-resource languages like English, they struggle with the same query in low-resource languages. These discrepancies emphasize a fundamental limitation in the understanding and reasoning capabilities of LLMs for low-resource languages.
To alleviate this issue, we propose a two-stage instruction tuning framework (LinguaLIFT) for enhancing low-resource language tasks. An additional language alignment layer is initially integrated into the LLM to adapt a pre-trained multilingual encoder. LinguaLIFT then improves multilingual alignment via code-switched fine-tuning, using the code-switched translation data generated from unsupervised multilingual alignment lexicons The second stage fine-tunes LLMs with English-only instruction data while freezing the language alignment layer, which enables LLMs to transfer the learned task-specific capabilities from English to low-resource languages. Additionally, we introduce a new benchmark named Multilingual Math World Problem (MMWP), spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages to evaluate multilingual reasoning tasks comprehensively. Experiments demonstrate that LinguaLIFT significantly outperforms several strong competitive methods on the MMWP and other widely used benchmarks, such as MGSM (Shi et al., 2023), MSVAMP (Chen et al., 2023), XNLI (Conneau et al., 2018) and X-CSQA (Lin et al., 2021).
2 Related Work
Leveraging Powerful Cross-lingual Transfer Capabilities of Multilingual Models.
Recent studies have shown that multilingual models possess strong cross-lingual transfer capabilities (Conneau et al., 2020; Xue et al., 2021; FitzGerald et al., 2023; Chirkova and Nikoulina, 2024; Shaham et al., 2024; Chen et al., 2024c), benefiting the low-resource language community by enabling task transfer across multiple languages after fine-tuning on high-resource languages (Kew et al., 2023; Pfeiffer et al., 2020; Reimers and Gurevych, 2020; Pan et al., 2021; Feng et al., 2022).
Unlike previous approaches focusing on multilingual instruction tuning (Singh et al., 2024; Üstün et al., 2024; Köpf et al., 2024; Zhu et al., 2024b, a; Li et al., 2023a; Wei et al., 2023; Ranaldi et al., 2023), LinguaLIFT incorporates a language alignment layer into LLMs to adapt a pre-trained multilingual encoder through code-switched fine-tuning. While freezing the language alignment layer and a pre-trained multilingual encoder, the LLM is fine-tuned with English-only instruction data to transfer task-specific capabilities from English to low-resource language tasks.
Improving Multilingual Mathematical Reasoning Tasks.
Recent efforts to improve multilingual mathematical reasoning for LLMs can be categorized into three ways: 1) Prompting close-source LLMs: Qin et al. (2023), Shi et al. (2023), and Huang et al. (2023) designed prompts for closed-source LLMs like ChatGPT, translating non-English contexts into English for reasoning. However, this approach is limited by translation quality and does not improve multilingual understanding or work well for open-source LLMs (Zhu et al., 2024b). 2) Instruction-tuning open-source LLMs: Chen et al. (2023), Chai et al. (2024), and Lai and Nissim (2024) adopted a translate-training method, translating English reasoning datasets into non-English and instruction-tuning LLMs. Zhu et al. (2024b) and She et al. (2024) proposed approaches to transfer mathematical reasoning capabilities from English to non-English. While these methods improve multilingual reasoning, they incur high translation costs and errors, making them impractical for low-resource languages. 3) Bridging existing skilled LLMs to multilingualism: Yoon et al. (2024) combines pre-trained multilingual models with skilled reasoning LLMs, but a performance gap persists between low-resource and high-resource languages. While Huang et al. (2024) shows significant improvement, it relies heavily on parallel and multilingual task data.
Different from existing studies, this paper proposes a novel two-stage instruction fine-tuning framework to enhance the reasoning capabilities of low-resource language tasks without relying on multilingual instruction data.
Figure 2: Overview of the proposed LinguaLIFT approach. Stage-I (Language-Align): A language alignment layer is introduced into the LLM to adapt the pre-trained multilingual encoder, thereby enhancing multilingual alignment through code-switched tuning. Stage-II (Task-Transfer): The LLM is fine-tuned on high-quality, English-only instruction data with the language alignment layer frozen, allowing LLM to transfer task-specific capabilities learned from English to low-resource languages.
3 Methodology
Figure 2 illustrates an overview of the proposed two-stage instruction tuning framework, LinguaLIFT. The core idea is to transfer task-specific capabilities learned from English to low-resource languages, leveraging the language alignment established by code-switched tuning. We then introduce the model architecture (§3.1), the language alignment stage (§3.2) and the task transfer stage (§3.3).
3.1 Model Architecture
Given a task-specific input $x$ with $l_{x}$ tokens, we use a multilingual pre-trained model to encode it into a moderately language-agnostic representation $X$, facilitating more effective cross-lingual understanding and transfer:
$X=Encoder(x),$ (1)
where $Encoder(\cdot)$ is the pre-trained multilingual encoder, and $X\in\mathbb{R}^{l_{x}\times d_{1}}$ is the final hidden state.
We introduce the language alignment layer, $Alignment(\cdot)$, to bridge the gap between the multilingual encoder and the LLM representation, implemented as a multi-layer perceptron (MLP). This layer maps $X$ to the input embedding space of the LLM:
$\hat{X}=Alignment(X),$ (2)
where $\hat{X}\in\mathbb{R}^{l_{x}\times d_{2}}$ is the alignment representation of $X$ on the input embedding space of the LLM.
To leverage the built-in capabilities of the LLM for instruction following, we embed the instruction context $q$ with $l_{q}$ tokens using the embedding layer $Embedding(\cdot)$:
$Q=Embedding(q),$ (3)
where $Q\in\mathbb{R}^{l_{q}\times d_{2}}$ presents the instruction context $q$ in the input embedding space of the LLM. Then, we concatenate the instruction context representation with the multilingual input representation:
[…]
3.2 Language Alignment
In this stage, we enhance the model’s multilingual alignment through code-switched tuning. This involves two key parts: building multilingual alignment lexicons and enhancing alignment via code-switched translation instruction tuning.
Building Multilingual Alignment Lexicons.
Given the scarcity of data for low-resource languages, we adopt an unsupervised word translation method MUSE (Lample et al., 2018) to construct the multilingual alignment lexicons without parallel data. Specifically, we extract a set of unique words $\mathcal{W}=\{w\}_{i}^{N}$ from English monolingual corpus $\mathcal{D}$, where $N$ is the number of unique words, excluding named entities and stop words. We collect the translations for each word $w\in\mathcal{W}$ by retrieving the top nearest neighbor by projecting the word embeddings into the target space. Further implementation details are provided in Appendix A.1.
Enhancing Alignment with Code-Switched Translation Instruction Tuning.
We leverage the built multilingual alignment lexicons to generate code-switched inputs. Formally, this tuning aims to predict the English translation response tokens $y$ given the instruction context $q$ and code-switched input tokens $x_{l}$ in language $l$. The input embeddings for this task are represented as $(Q,\hat{X_{l}})$ from Equation (4). The optimization objective is outlined as follows:
[…]
where $T$ represents the sequence length of the target output $y$, $y_{<i}$ denotes the tokens preceding the $i$-th token, and $y_{i}$ is the $i$-th token of $y$. The trainable parameters include $\theta$, corresponding to the language alignment layer and the boundary tokens. $\sigma$ and $\phi$ represent the parameters of the multilingual encoder and LLM, respectively, which remain frozen during training to mitigate the risk of catastrophic forgetting. Finally, $\mathcal{L}$ denotes the set of target languages.
3.3 Task Transfer
After establishing language alignment in the first stage, we fine-tune LLM using high-quality, English-only task-specific instruction data while freezing the language alignment layer. This approach preserves the multilingual alignment and cross-lingual transferability capabilities acquired in the first stage, unlocking the potential of LLM to transfer strong capabilities learned from English to low-resource language tasks. Specifically, we fine-tune the parameters of the LLM to minimize the language modeling negative log-likelihood of the response $y$, given the combined instruction context and task-specific input $(Q,\hat{X}_{en})$:
$-\underset{\phi}{\arg\min}\sum_{i=1}^{T}\log p_{\tilde{\theta},\sigma,\phi}\left(y_{i}\mid(Q,\hat{X}_{en}),y_{<i}\right),$ (6)
where $\phi$ represents the trainable parameters of the LLM, and $\tilde{\theta}$ is initialized from the checkpoint of $\theta$ trained at the first stage. The multilingual encoder parameters $\sigma$ and $\tilde{\theta}$ are kept frozen throughout the training.
4 The MMWP Benchmark
Existing multilingual mathematical reasoning benchmarks (Shi et al., 2023; Chen et al., 2023) predominantly focus on 7 high-resource languages and 3 low-resource languages, leaving a significant gap in coverage for low-resource languages. This imbalance introduces evaluation bias, as models are optimized for high-resource languages while their performance on low-resource languages remains underexplored, hindering the comprehensive development of multilingual models. To fill this gap, we build a new Multilingual Math World Problem (MMWP) benchmark and describe its collection process in this section.
Source data.
We use AsDiV (Miao et al., 2020) and MAWPS (Koncel-Kedziorski et al., 2016) as base datasets. From the AsDiv official test set, we randomly select 500 examples and another 500 from MAWPS, where all the problems require multiple steps to solve, as described by Miao et al. (2020). We filter out duplicates and problems with non-numeric answers, resulting in 811 examples.
Target Language Selection.
We select a typologically diverse set of 48 languages, spanning 12 language families and 12 writing systems, with a range of resource levels—low, medium, and high. This includes 21 low-resource languages [^2] Afrikaans (af), Arabic (ar), Belarusian (be), Bengali (bn), Basque (eu), Gujarati (gu), Hausa (ha), Hindi (hi), Armenian (hy), Icelandic (is), Kannada (kn), Luxembourgish (lb), Macedonian (mk), Malayalam (ml), Marathi (mr), Nepali (ne), Slovak (sk), Swahili (sw), Tamil (ta), Telugu (te), Thai (th), 17 medium-resource languages [^3] Bulgarian (bg), Catalan (ca), Czech (cs), Danish (da), Finnish (fi), Croatian (hr), Hungarian (hu), Indonesian (id), Korean (ko), Norwegian Bokmål (nb), Polish (pl), Portuguese (pt), Romanian (ro), Slovenian (sl), Serbian (sr), Ukrainian (uk), Vietnamese (vi), and 10 high-resource languages [^4] German (de), English (en), Spanish (es), French (fr), Italian (it), Japanese (ja), Dutch (nl), Russian (ru), Swedish (sv), Chinese (zh). Additional selection details are provided in Appendix B.1.
Translation process.
To ensure the high quality of translations, we employed Google Translation System to translate the selected English questions into 47 other languages. Moreover, five annotators then post-edit and calibrate the translation, followed by quality estimation through both human and automatic evaluation, as detailed in Appendix B.2.
5 Experiments
5.1 Datasets
Evaluation Dataset.
We use the MMWP and latest multilingual benchmarks, MGSM (Shi et al., 2023) and MSVAMP (Chen et al., 2023) to evaluate the performance of LLMs in multilingual mathematical reasoning. We evaluate in zero-shot chain-of-thought reasoning (Wei et al., 2022) setting. To further assess the task generalization of LinguaLIFT, we incorporate several challenging multilingual datasets, including X-CSQA (Lin et al., 2021) for commonsense reasoning and XNLI (Conneau et al., 2018) for natural language inference.
Training Dataset.
We utilize English-only instruction data following prior work (Lu et al., 2024; Huang et al., 2024), which include MetaMathQA (Yu et al., 2024) for mathematical reasoning, MultiNLI (Williams et al., 2018) for natural language inference, and a set of unified commonsense reasoning tasks, comprising the X-CSQA, OpenBookQA (Mihaylov et al., 2018), ARC (Clark et al., 2018), and QASC (Khot et al., 2020) datasets. Additionally, to further explore the potential of the proposed method, we incorporate the recent advanced mathematical reasoning instruction dataset, OpenMathInstruct-2 (Toshniwal et al., 2024). Statistics of the datasets involved are presented in Table 5, and the prompts for each task are given in Appendix A.3.
5.2 Baselines
We consider three categories of baselines: (1) Mono-SFT (Luo et al., 2023; Yue et al., 2024; Yu et al., 2024; Toshniwal et al., 2024; Zhu et al., 2024b), a vanilla method that fine-tunes the model on English task datasets. (2) Multi-SFT (Chen et al., 2023; Zhu et al., 2024b; She et al., 2024), a translation-based method that fine-tunes the model on multilingual task datasets. (3) Leveraging External Tools or Models (Shi et al., 2023; Yoon et al., 2024; Huang et al., 2024), a method that leverages external translation systems or models with existing powerful LLMs. More training details are presented in Appendix A.4.
5.3 Experimental Results
LLaMA-2-7B as base model LR. MR. HR. Avg.
baseline
Mono-SFT
MAmmoTH 6.36 20.4 26.1 15.5
WizardMath 10.1 28.0 32.2 21.0
MetaMath 13.6 37.2 41.8 27.8
OpenMath2 19.3 55.5 62.7 41.1
Multi-SFT
MathOctopus-Parallel 11.0 20.8 25.2 17.4
MathOctopus-MAPO-DPO 18.6 29.8 33.1 25.6
MetaMathOctopus-MAPO-DPO 16.4 36.7 43.9 29.3
QAlign-MetaMathQA 17.1 39.0 44.5 30.6
Leveraging External Tools or Models
Translate-En (MetaMath) 27.6 36.4 40.6 33.4
LangBridge (MetaMath) 33.4 36.8 39.4 35.9
MindMerger-Soft (MetaMath) 36.6 40.7 43.2 39.4
(Add Low-Resource) 36.7 39.8 41.5 38.8
(Low-Resource Retrain) 33.4 36.8 39.4 35.9
Translate-En (OpenMath2) 43.1 54.7 61.5 51.0
LangBridge (OpenMath2) 47.9 53.8 59.2 52.3
MindMerger-Soft (OpenMath2) 49.5 54.8 61.2 53.8
our methods
LinguaLIFT (MetaMath) 41.2 44.3 45.5 43.2
LinguaLIFT (OpenMath2) 55.4 61.2 62.5 58.9
Table 1: Results on the MMWP benchmark. "LR." denotes the average performance for low-resource languages, "MR." for medium-resource languages, and "HR." for high-resource languages. The bold text highlights the highest score among systems. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation. The LLM and extra data used in the third baseline category are indicated in brackets.
LinguaLIFT demonstrates significant improvements across low-resource languages on MMWP.
Experimental results on the MMWP test set, grouped by language resource levels, are presented in Table 1. We elaborate on six key observations from the results: (1) Mono-SFT models exhibit substantial performance drops in low-resource languages and slight drops in medium-resource languages. (2) Multi-SFT models, while outperforming their monolingual counterparts in low-resource settings, still show a considerable performance gap between low-resource and high-resource languages. (3) Models leveraging external translation systems or pre-trained multilingual models achieve limited generalization to unseen languages, and they still lag behind in performance compared to high-resource languages. (4) MindMerger, retrained using additional open-source parallel data in low-resource languages for a fair comparison, performs well in high-resource languages while hardly improving low-resource tasks, demonstrating its limited applicability in low-resource scenarios. (5) LinguaLIFT significantly enhances low-resource reasoning performance, outperforming all competitive baselines. (6) Incorporating advanced English reasoning datasets (e.g., OpenMathInstruct-2) further enhances low-resource language reasoning, highlighting the importance of adapting to evolving, high-quality data for improving low-resource language reasoning.
These results highlight the importance of comprehensively evaluating multilingual reasoning models across diverse languages and further validate LinguaLIFT’s effectiveness in enhancing reasoning performance for low-resource languages.
LinguaLIFT enhances LLM performance in low-resource languages on MGSM and MSVAMP test sets.
MGSM MSVAMP
LLaMA-2-7B as base model LR. HR. Avg. LR. HR. Avg.
baseline
Mono-SFT
MAmmoTH 3.40 32.4 21.9 6.57 40.1 30.1
WizardMath 4.00 37.7 25.5 15.7 48.5 38.7
MetaMath 4.60 51.7 34.6 15.2 61.2 47.4
OpenMath2 5.60 60.2 40.4 18.8 70.5 55.0
Multi-SFT
MathOctopus-Parallel 28.0 42.4 37.2 33.6 43.8 40.8
MathOctopus-MAPO-DPO 30.6 43.4 38.8 52.5 58.8 56.9
MetaMathOctopus-MAPO-DPO 31.0 55.6 46.7 57.8 67.4 64.5
QAlign-MetaMathQA 26.3 55.9 45.2 48.4 61.5 57.6
Leveraging External Tools or Models
Translate-En-MetaMath 39.6 55.1 49.4 47.4 52.3 50.9
LangBridge-MetaMath 38.4 51.9 47.0 43.9 54.5 51.4
MindMerger-Soft-MetaMath 53.1 57.9 56.2 52.7 60.6 58.2
Translate-En-OpenMath2 41.3 63.8 55.6 51.9 62.2 59.1
LangBridge-OpenMath2 42.6 60.8 54.2 47.9 64.4 59.5
MindMerger-Soft-OpenMath2 60.5 67.5 65.0 63.4 74.2 70.9
our methods
AlignIFT-MetaMath 55.4 58.8 57.6 56.1 60.6 59.3
AlignIFT-OpenMath2 63.8 66.5 65.5 67.2 74.3 72.2
Table 2: Results on MGSM and MSVAMP datasets. "LR" denotes the average multilingual performance for low-resource languages, while "HR" represents the performance for high-resource languages. The bold text highlights the highest score among systems. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation.
Experimental results on the other two widely used multilingual reasoning benchmarks MGSM and MSVAMP are presented in Table 2. These findings are generally consistent with those from the MMWP test sets. LinguaLIFT notably outperforms previous competitive baselines in low-resource languages on both MGSM and MSVAMP test sets, highlighting its strong generalization to various out-of-domain reasoning tasks and robust performance across low-resource language mathematical reasoning tasks.
Versatile Adaptation Across Various Tasks with LinguaLIFT.
XNLI X-CSQA
LLaMA-2-7B as base model LR. HR. Avg. LR. HR. Avg.
baselines
Mono-SFT(English-only Task Data)
Mono-SFT* 58.7 80.1 68.7 28.6 58.6 51.3
Multi-SFT(Multilingual Task Data with Query Translation)
Multi-SFT* 63.6 81.5 71.9 29.4 48.6 43.8
QAlign* 67.1 80.9 73.5 35.5 57.9 52.3
Freezing LLM with External Tools or Models
LangBridge* 73.9 79.4 76.5 30.9 37.8 36.1
Translate-En* 71.1 79.6 75.1 42.7 55.5 52.3
MindMerger-Soft* 74.4 83.1 78.4 47.9 65.4 61.0
our methods
LinguaLIFT 77.6 83.3 80.3 49.6 65.5 61.5
Table 3: Results on XNLI and X-CSQA datasets. "LR" denotes the average performance for low-resource languages, and "HR" denotes high-resource languages. The bold text highlights the highest score among systems. The asterisk symbol (*) indicates results are taken from the published results of Huang et al. (2024).
As shown in Table 3, LinguaLIFT outperforms all baselines on language understanding tasks, such as XNLI and X-CSQA, demonstrating its efficiency in adapting various tasks than previous work, as it only needs English-only instruction data while others need to translate instruction data to target language constantly. These results highlight LinguaLIFT’s versatility in generalizing beyond mathematical reasoning, making it a universal and effective approach for enhancing low-resource language task performance.
LinguaLIFT Generalizes well on High-Resource Tasks.
Although LinguaLIFT is designed for low-resource languages, it also boosts performance in high-resource languages. LinguaLIFT outperforms several competitive baselines on mathematical reasoning tasks and multilingual understanding tasks, as shown in Tables 1, 2 and 3. These results highlight LinguaLIFT’s ability to improve both reasoning and comprehension across low- and high-resource languages.
(a) Tow-stage Training
(b) Trainable Modules
Figure 3: Ablation experiments of two-stage training and trainable modules on MGSM. LR., HR., and Avg. represent the average accuracy across low-resource languages, high-resource languages, and all languages, respectively.
Overall, LinguaLIFT demonstrates its effectiveness in enhancing low-resource language capabilities without relying on multilingual instruction data. It established itself as a universally applicable and strong potential method for LLMs to advance low-resource language performance across various tasks. More experimental results are provided in Appendix C.
6 Analysis
6.1 Two-Stage Training Ablation Studies
Training Stage Ablation
As shown in Figure 3a, removing the Language-Align stage leads to a noticeable performance drop in low-resource languages, demonstrating the essential of the Language-Align stage in enhancing low-resource language tasks. Removing the Task-Transfer stage causes substantial performance degradation across all languages. These results suggest the necessity of the two stages. Further details are provided in Appendix D.1.
Figure 4: Ablation experiments on the impact of varying code-switch ratios on multilingual alignment and reasoning performance.
Trainable Modules Ablation
As illustrated in Figure 3b, training the LLM during the Language-Align stage reduces performance on high-resource language reasoning. Freezing LLM training in the Task-Transfer stage significantly degrades reasoning performance across all languages. Additionally, training the language alignment layer during the Task-Transfer stage harms performance on low-resource language reasoning. These results indicate the benefit of first training the language alignment layer, followed by LLM fine-tuning. Further details are available in Appendix D.2.
6.2 Analysis of Code-Switched Tuning
Impact of Code-Switched Ratio to Multilingual Alignment and Reasoning Tasks.
Figure 5: T-SNE visualization in the spaces of the Mono-SFT input embeddings, MindMerger mapping layer outputs, and LinguaLIFT language alignment layer outputs.
We report top-1 retrieval accuracy and reasoning accuracy on Tatoeba (Artetxe and Schwenk, 2019) and MGSM (Shi et al., 2023) to indicate language alignment degree and reasoning performance, respectively, as shown in Figure 4. Multilingual alignment improves as the code-switch ratio increases, especially for low-resource languages, and reasoning performance also increases. Notably, at an 80% code-switch ratio, the model’s reasoning performance reaches comparable performance to those trained with parallel corpora.
Impact of the Part-of-Speech in Code-Switched to Reasoning tasks.
Figure 6: The reasoning performance on MGSM group by the combinations of POS: (1) individual POS categories:Verb (V), Adposition (ADP), Pronoun (Pron); (2) syntactic function combinations:Verb+Adverb (V+Adv), Adjective+Adverb (Adj+Adv), Pronoun+Auxiliary+Conjunction (Pron + Aux + Conj); and (3) key syntactic structures:Subject-Verb (N+V) and Prepositional Phrases (N+ADP).
As shown in Figure 6, among the individual part-of-speech (POS) groups, substituting nouns significantly affects the model’s reasoning ability. Syntactic structures involving subject-verb and prepositional phrase combinations outperform other POS combinations, highlighting the importance of core arguments (subjects and verbs) and their relations (prepositions) for capturing key relationships in reasoning tasks. These structures are sufficient for the model to generalize reasoning across languages. In contrast, adjective-adverb and auxiliary-conjunction combinations perform the worst, suggesting that modifiers are less critical for reasoning tasks. More details are provided in Appendix D.3
6.3 LinguaLIFT better aligns low-resource languages
For each low-resource language in the MMWP benchmark, we selected 100 texts with equivalent meaning from the Flores-101 dataset (Goyal et al., 2022). We obtained these input mean pooling representations from different methods and visualized them using T-SNE (van der Maaten and Hinton, 2008). As shown in Figure 5, the LLM embeddings of low-resource languages are distinct from English, indicating the challenges of understanding and transferring knowledge to these languages. In the case of MindMerger, some low-resource language representations closely overlap with English, while others remain isolated, demonstrating its limitations in scenarios involving a wide range of low-resource languages. In contrast, LinguaLIFT aligns low-resource language representations more closely with English, enabling better transfer of reasoning capabilities from English instruction data, thereby resulting in improved reasoning performance.
6.4 Supplementary Experiments
We conducted several supplementary experiments, including the quantitative analysis of the correlation between multilingual alignment and reasoning performance (Appendix E.1), the analysis of language transferability in language families and writing systems (Appendix E.2), the adaptation of different types and scales of LLMs (Appendix E.3, the selection of the language alignment layer and pre-trained multilingual encoder (Appendix E.4, E.5), the incorporation of multilingual instruction data into LinguaLIFT (Appendix E.6), and zero-shot CoT examples in mathematical reasoning tasks (Appendix E.7).
7 Conclusion
This paper introduced LinguaLIFT, a novel two-stage instruction tuning framework that enhances low-resource language tasks without relying on parallel corpora or multilingual instruction data. Additionally, we introduced MMWP, a multilingual benchmark spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages, to comprehensively evaluate multilingual mathematical reasoning tasks. Experiments on the MMWP and other widely used benchmarks demonstrate its effectiveness in advancing low-resource language tasks and further alleviating the performance gap between high-resource and low-resource language tasks in LLMs.
Limitations
While our experimental results demonstrate that the proposed two-stage instruction tuning method significantly improves low-resource language reasoning and understanding tasks, it does require a certain level of computational resources. Specifically, the need for a moderately larger pre-trained multilingual encoder and full fine-tuning of the LLM may impose a computational burden. Additionally, we have not explored parameter-efficient fine-tuning (PEFT) methods, so the effectiveness of our approach in PEFT settings remains untested. Future work could focus on developing more efficient and lightweight tuning strategies for LinguaLIFT to reduce the computational limits.
Appendix A Implementation Details
A.1 Collection of the Alignment Lexicons
To better construct multilingual alignment lexicons for low-resource language tasks, we leveraged the Unsupervised Bilingual Lexicon Induction (UBLI) (Zhang et al., 2017; Lample et al., 2018; Dou et al., 2018; Artetxe et al., 2019; Li et al., 2023b), which has been proven effective in inducing word translation pairs by aligning independently trained word embeddings in two languages.
Initially, we tokenized the text using the Spacy library’s tokenization function [^5] https://spacy.io/. The resultant word set comprised all words, barring specific named entities, numbers, and date tokens. This step is crucial in ensuring our focus on frequent and pertinent terms that are likely to hold significance in task-specific domains.
Following prior outstanding work MUSE [^6] https://github.com/facebookresearch/MUSE (Lample et al., 2018), we construct multilingual lexicons using adversarial training to establish a linear mapping between source and target spaces without relying on cross-lingual supervision. This process involves training the model to align the word embeddings of the source and target languages in a shared semantic space. For each word in the source language, we identify the most relevant translations in the target language by projecting the source word embeddings into the target space and retrieving the top nearest neighbor words based on cosine similarity. These translated words form the bilingual lexicons, which are essential for enhancing multilingual understanding. Table 4 provides examples of word translations derived from the multilingual alignment lexicons we constructed, illustrating the effectiveness of this approach.
Table 4: Word translations examples for English to several low-resource languages.
This unsupervised approach allowed us to obtain diverse translations while preserving the general meaning of the words, which is crucial for capturing cross-lingual alignment. Moreover, by avoiding context-specific translations, we ensured that the generated translations had broad applicability across multiple tasks and languages. This strategy reduced the reliance on large-scale parallel corpora or pre-existing bilingual dictionaries and provided a robust and scalable solution for building multilingual lexicons in an unsupervised manner.
A.2 Dataset Statistics
Dataset # Lang # Samples
Traing Datasets
MetaMathQA 1 395,000
OpenMathInstruct-2 1 1,000,000
MGSM8KInstruct 10 73,599
MetaMathQA 1 395,000
Multi-NLI 1 392,702
QASC 1 8,134
ARC 1 3,370
OpenBookQA 1 4,957
X-CSQA(En) 1 8,888
Evaluation Datasets
MMWP 48 38,400
MGSM 11 2,750
MSVAMP 10 10,000
XNLI 15 75,150
X-CSQA 16 16,000
Table 5: Statistics of involved datasets. "#Langs" denotes the number of languages covered by the dataset, and "#Sample" refers to the total number of samples it contains. "Scenario" denotes the usage scenario of the specific dataset discussed in the methodology section.
A.3 Instruction Tuning Prompts
The prompt for code-switched tuning is adapted from Zhang et al. (2023), where the source language, source sentence, and target language are replaced with the relevant translations.
The prompts for mathematical reasoning tasks, natural language inference, and commonsense question answering are modified from Toshniwal et al. (2024) and Lu et al. (2024), with the general instruction being replaced by the specific problems from the training data.
Prompt for Code-Switched Tuning Translate the following code-switched sentence from {source_lang} to pure {target_lang}: {source_lang}: {source_sentence}
{target_lang}
[…]
Prompt for XNLI I will give you a premise and a hypothesis. Choose the most appropriate relationship from the following options: Entailment, Neutral, Contradiction. ### Premise: {premise} ### Hypothesis: {hypo} ### Answer:
Prompt for X-CSQA Question: {question}
Choices: A. {choice_A} B. {choice_B} C. {choice_C} D. {choice_D} E. {choice_E}
Answer:
A.4 Training Details
We use LlamaFactory [^7] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024) as the training codebase for our experiments. In the first stage, we train only the additional alignment modules (e.g., the language alignment layer and the boundary tokens) for 3 epochs, with a constant learning rate of 6e-4 and a batch size of 256. In the second stage, we fine-tune all parameters of the LLM for 3 epochs. The learning rate is set to 2e-5, with a warm-up ratio of 0.05 and a cosine learning rate scheduler. We also apply a weight decay of 1e-2 and use a batch size of 128. All experiments are conducted on eight NVIDIA A100 GPUs for a day.
Appendix B Details of Constructing the MMWP Benchmark
B.1 Selection of Target Languages and Categorization of Resource-level.
In our study, we employ a unique categorization method for selecting target languages, which deviates somewhat from the conventional definitions based on the abundance or scarcity of linguistic resources. This deviation is primarily due to our focus on LLMs. Rather than adhering to traditional classifications, we opt to categorize languages based on their language distribution in the pre-training corpus used for the LLMs.
Guided by the LLaMA2 technical report (Touvron et al., 2023), we define low-resource languages as those that constitute less than 0.005% of the available multilingual datasets. On the other hand, we categorize languages with a representation percentage between 0.005% and 0.1% as medium-resource languages. This categorization method allows us to incorporate a diverse set of languages into our study, thereby enabling a more effective and wide-ranging assessment of the multilingual reasoning capabilities of our model.
Furthermore, we believe that the comprehensive evaluation of multilingual tasks necessitates the inclusion of languages from various families and scripts. This diversity is crucial in understanding the robustness and versatility of our model, as it allows us to evaluate its performance and generalization capabilities across different linguistic contexts and resource levels. By incorporating languages from various families, scripts, and resource categories, we can ensure more comprehensive coverage in our multilingual settings evaluation.
This unique approach to language selection and categorization provides a more nuanced understanding of language resource levels in LLMs while also ensuring a broader and more diverse evaluation of multilingual tasks for LLMs. It also underscores the importance of considering language families and scripts for the assessment of LLMs, thereby contributing to a more comprehensive and inclusive approach to language model development and evaluation.
B.2 Quality Estimation proves the effectiveness of the proposed MMWP benchmark.
To ensure the quality of the proposed MMWP benchmark, we conducted both human and automatic quality estimation evaluations.
For human evaluation, we employed five annotators, each of whom was responsible for post-calibrating and assessing the quality of the translated dataset. To facilitate this, we utilized GPT-4 for back-translation of the problems into English, a process that allowed the annotators to compare the translated versions with their original counterparts. The evaluators were tasked to scrutinize the translations for grammatical correctness, fluency, and semantic accuracy. Any significant deviations in meaning between the original and translated versions were carefully addressed. The annotators were guided to revise such instances to better align with the original intent, thereby ensuring the preservation of the original linguistic integrity. Moreover, the post-editing process was carried out to maintain native language properties, such as syntactical flexibility and natural phrasing, thus ensuring the fluency and native-like quality of the back-translations. This rigorous human calibration process aimed to certify that the final MMWP dataset was linguistically accurate and culturally appropriate for all represented languages.
The prompt for back-translating the low-resource language texts and assisting the post-editing process are presented below.
GPT-4 Prompt for Back-Translating You are a translation assistant. Directly translate the mathematical problems from {source_lang} to English without additional explanations. {source_sentence}
GPT-4 Prompt for Assisting Annotators Post-Editing You are provided with the following: 1. A low-resource source language text. 2. The back-translated English text (BT). 3. The reference English translation (REF). Your task is to evaluate the quality of the back-translated English text (BT) based on three criteria: grammatical correctness, fluency, and semantic accuracy. Then, propose three alternative revisions to improve the BT. For each revision, explain why it was made and how it improves the translation. 1. Grammatical Correctness:
Does the back-translated text adhere to standard English grammar rules (e.g., subject-verb agreement, punctuation, tense consistency)?
2. Fluency:
Is the back-translated text natural and smooth? Does it sound like it was written by a native speaker?
3. Semantic Accuracy:
Does the back-translated text accurately reflect the meaning of the source language text? Are there any discrepancies in the interpretation of the source?
Provide three revision suggestions for improving the BT: 1. Each revision should aim to enhance either grammatical correctness, fluency, or semantic accuracy. 2. For each suggestion, explain: Why the revision is necessary. Which aspect of translation quality (grammar, fluency, or accuracy) it improves.
[…]
Average 78.71 86.08 1.09
Table 6: Automatic quality estimation of MMWP using back-translations, showing average results across all languages.
Complementing the human evaluation, we conducted an automatic evaluation to assess the MMWP dataset’s translation quality quantitatively. In this process, we adopted a back-translation strategy, translating the MMWP problems into the target languages and then back into English. Subsequently, we used widely-accepted automatic evaluation metrics such as BLEU (Papineni et al., 2002), chrF (Popović, 2015), and TER (Olive et al., 2011) to compare the back-translated problems with the original MMWP dataset. These metrics provided a quantitative measure of the overlap between the original and back-translated versions, thereby offering a quantifiable estimation of the translation fidelity. The results of this quality estimation, presented in Table 6, attest to the reliable quality of the translations.
The results from both the human and automatic evaluations indicate that the dataset is properly constructed and adequately reflects the multilingual nature of the tasks. The human post-editing process ensured that the translated problems maintained their semantic integrity, while the automatic evaluation confirmed that the translations preserved the necessary linguistic structure for multilingual reasoning. Overall, the quality assurance measures we implemented guarantee that the MMWP benchmark is reliable and effective in evaluating multilingual mathematical reasoning across diverse languages and resource levels.
Appendix C Complete Experimental Results
C.1 Evaluation Results on MMWP
The complete experimental results on MMWP are shown in Table 7, 8 and 9. These tables present Mono-SFT, Multi-SFT, and Leveraging External Tools or Models category comparison baselines.
Resource Level Language MAmmoTH-7B WizardMath-7B MetaMath-7B OpenMath2
Low-Resource af 20.0 25.3 32.4 48.1
ar 14.8 18.6 23.6 33.1
be 7.6 12.5 17.9 30.3
bn 4.8 7.9 9.9 15.7
eu 2.2 4.2 6.3 6.5
gu 2.5 3.6 4.8 4.0
ha 2.8 4.4 5.9 6.4
hi 7.6 17.0 24.5 34.4
hy 3.2 3.7 5.8 6.2
is 6.5 10.1 14.3 24.3
kn 2.1 3.6 6.4 4.8
lb 6.8 10.9 14.6 25.7
mk 15.9 20.1 30.1 45.6
ml 2.5 4.2 5.2 5.7
mr 3.0 8.4 9.6 14.9
ne 2.0 5.8 8.4 12.3
sk 15.0 23.7 32.3 47.5
sw 3.1 6.5 7.5 8.4
ta 2.6 6.0 6.3 5.4
te 2.5 4.9 6.2 4.4
th 6.0 11.6 13.4 21.5
Average 6.4 10.1 13.6 19.3
Medium-Resource bg 23.2 29.1 38.7 59.7
ca 23.9 31.4 38.8 59.1
cs 19.7 28.7 37.5 56.7
da 23.4 30.3 40.3 55.7
fi 17.3 22.9 32.6 51.1
hr 19.9 24.8 35.1 54.5
hu 13.4 24.3 31.0 47.2
id 22.2 30.0 38.4 57.6
ko 14.4 25.9 35.0 47.4
nb 23.2 29.4 37.9 56.5
pl 21.2 28.0 37.9 57.3
pt 25.5 36.3 42.9 65.0
ro 20.4 27.1 39.5 56.1
sl 18.3 24.5 33.1 51.7
sr 18.4 26.8 36.1 54.0
uk 20.7 27.6 40.3 58.8
vi 21.8 28.4 37.0 54.8
Average 20.4 28.0 37.2 55.5
High-resource de 25.2 32.3 41.9 62.8
en 36.6 37.6 42.7 75.0
es 30.2 33.8 43.4 66.7
fr 28.7 30.8 43.5 64.4
it 28.1 34.7 42.9 63.1
ja 19.7 28.6 40.8 56.4
nl 23.9 31.2 41.2 58.9
ru 23.9 30.3 41.6 60.8
sv 24.2 30.3 42.5 59.7
zh 20.7 32.1 37.9 59.2
Average 26.1 32.2 41.8 62.7
Table 7: Detailed results(Accuracy) of Mono-SFT baselines on the MMWP benchmark across 48 languages.
Resource Level Language MathOctopus-Parallel-7B MathOctopus-MAPO-DPO-7B MetaMathOctopus-MAPO-DPO-7B QAlign-MetaMathQA-7B
Low-Resource af 19.1 27.5 32.2 37.4
ar 8.6 17.1 20.4 22.4
be 14.6 23.2 21.8 18.1
bn 20.1 28.1 28.4 26.6
eu 5.6 11.2 9.4 7.5
gu 6.2 10.9 5.3 6.3
ha 5.6 10.6 8.8 6.3
hi 11.1 24.3 14.7 22.2
hy 6.0 8.4 6.0 6.5
is 10.5 18.9 15.7 17.4
kn 5.4 13.9 5.2 5.2
lb 8.8 15.5 18.0 14.7
mk 15.9 26.5 32.2 31.3
ml 5.9 9.9 7.3 6.7
mr 6.8 15.2 5.4 11.0
ne 6.5 16.3 9.1 10.2
sk 16.3 25.8 33.2 31.6
sw 21.7 26.9 30.7 30.2
ta 4.0 11.0 4.1 6.8
te 10.2 17.9 4.0 5.1
th 21.3 31.1 33.8 35.4
Average 11.0 18.6 16.4 17.1
Medium-Resource bg 20.6 30.1 36.4 39.2
ca 22.8 32.9 41.8 41.2
cs 18.9 28.0 36.3 39.1
da 22.1 28.7 39.5 40.7
fi 19.2 27.1 35.5 36.7
hr 21.5 28.7 33.1 36.4
hu 19.1 30.0 31.7 35.4
id 20.8 29.1 36.5 39.5
ko 18.0 29.4 33.4 38.5
nb 22.6 28.5 40.3 40.3
pl 21.0 31.3 38.1 39.2
pt 24.3 32.4 44.1 43.8
ro 19.6 30.7 36.4 38.8
sl 19.2 30.2 32.6 36.3
sr 18.5 29.4 33.3 37.2
uk 24.8 31.2 37.7 41.1
vi 20.4 28.9 36.6 39.3
Average 20.8 29.8 36.7 39.0
High-resource de 27.7 33.5 41.3 45.1
en 29.6 36.0 53.5 52.9
es 25.8 33.7 46.4 46.0
fr 24.8 32.4 44.1 43.4
it 24.0 31.6 44.8 43.4
ja 25.4 33.5 40.6 43.9
nl 21.7 30.7 39.6 42.2
ru 26.5 33.8 43.5 44.1
sv 21.1 30.5 40.9 42.2
zh 25.8 35.0 44.8 41.7
Average 25.2 33.1 43.9 44.5
Table 8: Detailed results of Multi-SFT on the MMWP benchmark across 48 languages.
Resource Level Language Translate-En Langbridge MindMerger AlignIFT-MetaMath
Low-Resource af 33.0 40.7 40.2 46.4
ar 34.9 33.4 35.8 42.5
be 31.2 35.1 37.1 39.6
bn 27.7 31.9 37.6 39.5
eu 18.8 30.2 35.4 39.6
gu 21.6 31.8 35.8 41.0
ha 19.9 28.0 29.1 34.4
hi 27.4 35.3 39.4 41.7
hy 31.5 33.3 34.7 39.7
is 31.5 34.7 35.6 41.9
kn 26.7 29.7 33.8 37.4
lb 29.2 36.0 37.6 43.4
mk 35.9 40.8 40.7 43.9
ml 19.3 28.4 37.4 40.7
mr 25.6 30.5 34.3 37.6
ne 25.9 31.8 37.9 40.6
sk 29.2 36.4 38.6 42.5
sw 27.3 41.2 36.4 47.6
ta 24.4 27.1 35.0 40.5
te 28.5 29.7 37.1 41.9
th 30.3 34.9 38.7 41.7
Average 27.6 33.4 36.6 41.2
Medium-Resource bg 38.4 40.3 40.6 46.5
ca 35.3 37.0 41.7 45.4
cs 37.7 37.4 44.1 47.4
da 39.4 39.5 42.1 46.0
fi 34.1 33.4 37.7 41.3
hr 31.0 33.7 37.4 40.2
hu 36.1 33.9 37.1 41.4
id 38.0 36.9 39.8 43.0
ko 30.7 33.1 39.5 42.6
nb 38.2 40.0 40.9 48.1
pl 38.8 38.5 40.6 43.8
pt 42.4 41.7 42.9 49.1
ro 38.0 37.5 42.3 43.7
sl 34.4 34.8 40.4 44.1
sr 38.1 36.7 42.4 44.8
uk 36.3 37.9 41.9 43.7
vi 31.6 33.9 39.8 42.0
Average 36.4 36.8 40.7 44.3
High-resource de 41.0 40.7 43.9 44.6
en 42.5 47.1 47.8 49.7
es 40.6 39.6 43.9 49.0
fr 40.2 39.3 42.5 45.6
it 42.6 41.9 42.7 47.7
ja 39.6 37.0 40.2 40.4
nl 39.8 41.9 43.5 47.5
ru 38.9 40.4 42.4 46.5
sv 38.8 31.0 42.2 41.4
zh 41.5 35.1 43.3 42.1
Average 40.6 39.4 43.2 45.5
Table 9: Detailed results(Accuracy) of Leveraging External Tools or Models baseline and LinguaLIFT on the MMWP benchmark across 48 languages.
C.2 Evaluation Results on MGSM and MMWP
The complete experimental results on MGSM and MSVAMP are shown in Table 10, and 11.
LLaMA-2-7B as base model Bn De En Es Fr Ja Ru Sw Te Th Zh LR. HR. Avg.
Baseline
Mono-SFT
MAmmoTH 3.20 33.2 47.2 34.4 32.4 23.2 30.4 2.00 1.60 6.80 26.0 3.40 32.4 21.9
WizardMath 4.40 37.2 51.6 43.2 39.2 27.2 36.8 4.40 1.60 5.60 28.8 4.00 37.7 25.5
MetaMath 5.60 56.4 68.0 54.4 56.0 34.8 52.8 5.20 2.00 5.60 39.6 4.60 51.7 34.6
OpenMath2 6.80 65.6 79.2 66.0 62.4 42.4 57.6 6.00 2.80 6.80 48.4 5.60 60.2 40.4
Multi-SFT
MathOctopus-Parallel 31.2 46.0 51.6 42.8 43.6 34.0 40.0 36.4 10.8 33.6 38.8 28.0 42.4 37.2
MathOctopus-MAPO-DPO 33.2 47.2 46.8 43.2 39.6 41.6 40.8 37.2 13.6 38.4 44.8 30.6 43.4 38.8
MetaMathOctopus-MAPO-DPO 35.6 52.4 70.0 58.0 51.2 46.4 56.0 42.4 2.00 44.0 55.2 31.0 55.6 46.7
QAlign-MetaMathQA 29.6 54.0 68.4 57.6 59.2 45.2 58.4 35.6 2.40 37.6 48.8 26.3 55.9 45.2
Leveraging External Tools or Models
Translate-En-MetaMath 49.0 59.6 65.6 59.8 56.2 49.0 48.4 37.4 34.6 37.2 47.0 39.6 55.1 49.4
LangBridge-MetaMath 41.2 53.2 62.4 58.0 51.6 39.6 55.2 39.6 28.0 44.8 43.2 38.4 51.9 47.0
MindMerger-Soft-MetaMath 50.4 59.6 67.2 58.4 55.6 50.0 61.6 55.2 52.8 54.2 53.2 53.1 57.9 56.2
Translate-En-OpenMath2 52.1 69.8 76.9 71.6 63.4 57.4 54.2 38.4 35.6 39.2 53.2 41.3 63.8 55.6
LangBridge-OpenMath2 43.2 63.0 73.8 69.8 58.9 47.6 60.2 41.0 39.8 46.2 52.2 42.6 60.8 54.2
MindMerger-Soft-OpenMath2 60.0 69.4 77.8 74.2 63.4 57.8 71.8 61.2 57.2 63.6 58.4 60.5 67.5 65.0
Our Methods
LinguaLIFT-MetaMath 54.4 62.0 64.8 63.6 56.8 50.0 60.4 55.6 54.0 57.6 54.0 55.4 58.8 57.6
LinguaLIFT-OpenMath2 63.0 69.6 76.0 75.2 62.8 52.0 71.6 64.2 61.2 66.6 58.0 63.8 66.5 65.5
Table 10: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Yoon et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation.
LLaMA-2-7B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baselines
Mono-SFT
MAmmoTH 6.2 44.1 39.5 45.4 42.3 34.1 38.1 5.1 8.4 37.5 6.57 40.1 30.1
WizardMath 16.4 49.1 56.1 50.5 50.8 45.4 44.8 13.4 17.2 43.1 15.7 48.5 38.7
MetaMath 12.5 63.5 67.2 64.7 64.9 54.2 58.2 16.7 16.5 55.5 15.2 61.2 47.4
OpenMath2 19.9 72.7 78.6 72.5 72.9 64.6 68.5 13.1 23.5 63.8 18.8 70.5 55.0
Multi-SFT
MathOctopus-Parallel 27.8 43.8 46.1 45.4 44.9 41.4 43.8 38.1 34.9 41.4 33.6 43.8 40.8
MathOctopus-MAPO-DPO 48.3 57.6 59.0 59.4 60.0 58.7 58.0 54.6 54.7 59.1 52.5 58.8 56.9
MetaMathOctopus-MAPO-DPO 50.3 67.7 71.5 68.7 67.5 65.4 64.7 61.3 61.7 66.0 57.8 67.4 64.5
QAlign-MetaMathQA 41.9 63.5 65.5 63.0 62.8 58.0 61.5 53.6 49.7 56.2 48.4 61.5 57.6
Leveraging External Tools or Models
Translate-En-MetaMath 47.9 43.9 60.6 51.4 50.9 50.4 53.4 43.1 51.3 55.8 47.4 52.3 50.9
LangBridge-MetaMath 39.6 58.8 60.1 56.8 57.9 45.2 53.6 45.8 46.3 49.4 43.9 54.5 51.4
MindMerger-Soft-MetaMath 52.0 61.1 64.5 62.9 60.8 59.0 58.6 54.0 52.1 57.3 52.7 60.6 58.2
Translate-En-OpenMath2 56.3 54.1 72.0 59.5 59.3 60.9 64.6 40.2 59.2 64.9 51.9 62.2 59.1
LangBridge-OpenMath2 48.0 68.6 71.9 64.8 66.6 56.5 64.4 42.3 53.5 58.3 47.9 64.4 59.5
MindMerger-Soft-OpenMath2 61.2 77.2 77.1 73.6 73.5 74.1 70.8 64.8 64.1 72.8 63.4 74.2 70.9
our methods
LinguaLIFT-MetaMath 54.3 62.8 65.2 59.9 62.7 57.9 59.7 56.8 57.2 56.0 56.1 60.6 59.3
LinguaLIFT-OpenMath2 65.2 76.4 77.9 74.2 74.5 73.5 71.4 68.6 67.9 72.2 67.2 74.3 72.2
Table 11: Experimental Results on the MSVAMP Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Yoon et al., 2024), we classify Bn, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation.
C.3 Evaluation Results on XNLI and X-CSQA
The complete experimental results on XNLI and X-CSQA are shown in Table 10, and 11.
LLaMA-2-7B as base model Ar Bg Sw Th Tr Ur El Hi Zh Ru Vi De Fr Es En LR. HR. Avg.
baselines
Mono-SFT
Mono-SFT* 60.9 76.7 45.9 55.4 61.9 49.2 63.7 55.7 74.7 77.6 73.7 80.6 82.2 82.2 90.0 58.7 80.1 68.7
Multi-SFT
Multi-SFT* 61.7 78.7 56.3 60.1 65.6 57.5 67.0 61.7 79.1 79.7 73.7 82.3 82.9 83.9 88.8 63.6 81.5 71.9
QAlign* 67.0 79.4 65.2 65.2 67.9 62.2 66.5 63.3 76.6 79.2 73.7 80.9 83.1 83.8 89.1 67.1 80.9 73.5
Leveraging External Tools or Models
LangBridge* 75.2 79.6 71.7 72.4 74.8 66.9 79.1 71.1 77.4 77.4 78.5 78.8 79.9 80.5 83.4 73.9 79.4 76.5
Translate-En* 68.9 80.8 65.3 69.5 74.5 61.6 79.3 68.7 74.8 76 76.7 80.6 80.4 81.4 87.4 71.1 79.6 75.1
MindMerger-Soft* 76.2 82.4 66.6 71.8 75.7 69.4 78.5 74.7 80.0 80.7 80.3 83.5 83.9 84.4 88.7 74.4 83.1 78.4
our methods
LinguaLIFT 78.4 83.5 75.1 77.0 77.1 72.4 82.2 75.1 80.2 80.5 80.8 83.6 84.0 84.4 89.5 77.6 83.3 80.3
Table 12: Experimental Results on the XNLI Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. The asterisk symbol (*) indicates results obtained directly from the published results (Huang et al., 2024).
LLaMA-2-7B as base model Ar De En Es Fr Hi It Ja Nl Pl Pt Ru Sw Ur Vi Zh LR. HR. Avg.
baselines
Mono-SFT
Mono-SFT* 32.3 61.2 76.3 64.0 63.5 32.9 56.0 49.1 57.5 50.6 61.7 56.0 24.2 25.1 50.9 56.5 28.6 58.6 51.3
Multi-SFT
Multi-SFT* 28.7 49.1 67.2 54.3 52.1 32.0 50.2 38.7 45.9 45.5 51.2 46.5 27.6 29.2 38.8 43.8 29.4 48.6 43.8
QAlign* 36.3 58.8 75.7 63.1 60.3 37.8 58.3 49.2 56.3 51.3 59.8 56.3 35.1 32.6 50.5 54.8 35.5 57.9 52.3
Leveraging External Tools or Models
LangBridge* 30.6 37.4 44.4 38.4 38.2 30.6 39.1 33.9 38.4 39.8 36.3 35.1 31.8 30.5 33.3 39.8 30.9 37.8 36.1
Translate-En* 44.6 57.3 71.3 55.5 57.2 48.4 56.3 47.1 55.0 53.3 54.7 54.4 36.5 41.3 51.8 51.5 42.7 55.5 52.3
MindMerger-Soft* 51.4 67.0 78.1 69.1 68.1 48.4 66.8 53.9 63.8 63.3 67.1 63.7 45.5 46.2 60.6 62.9 47.9 65.4 61.0
our methods
LinguaLIFT 54.1 68.7 78.3 67.9 68.7 50.6 65.7 53.8 63.5 62.9 68.2 63.5 46.7 46.8 60.4 64.1 49.6 65.5 61.5
Table 13: Experimental Results on the X-CSQA Dataset. "LR." "HR." and "Avg." represent the average performance for details of the low-resource, high-resource, and all languages, respectively. The asterisk symbol (*) indicates results obtained directly from the published results (Huang et al., 2024).
Appendix D Analysis Experiments Details
This section outlines the experimental implementation for the various analysis experiments presented in the main text.
D.1 Ablation of Two-Stage Training
To demonstrate the necessity of the proposed two-stage instruction tuning approach, we perform an ablation study on the distinct components of the two-stage process: the Language-Align stage and the Task-Transfer stage. This study is conducted on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of each stage’s removal on reasoning performance across all languages. The detailed results are presented in Tables 14 and 15 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of the specific stage.
When the Language-Align stage was removed, we fine-tuned both the language-alignment layer and the LLM using only English reasoning data. This resulted in a noticeable decline in performance in low-resource languages, such as a 10.2% performance drop in Bengali (Bn), 13.2% performance drop in Thai (Th), 21.2% performance drop in Telugu (Te), and 7.6% performance drop in Swahili (Sw) on the MGSM dataset and a 14.4% performance drop in Bengali (Bn), 15.7% performance drop in Swahili (Sw) and 19.8% performance drop in Thai (Th) on the MSVAMP dataset. This decline highlights the importance of using code-switched multilingual input as a warm-up strategy before directly applying English-only instruction data. By incorporating this stage, LinguaLIFT can better leverage the multilingual model’s representation space, which contains valuable information from low-resource languages.
When the Task-Transfer stage was removed, we fine-tuned the language alignment layer using only code-switched translation data without updating the LLM’s parameters. After ablating the Task-Transfer stage, the models couldn’t complete reasoning tasks across all languages, resulting in an average accuracy of just 8.49% on the MGSM dataset and 9.74% on the MSVAMP dataset. This demonstrates that task-specific instruction tuning is critical for enabling the LLM to perform task-solving operations effectively.
These results underscore the complementary nature of the two stages: the Language-Align stage enhances the cross-lingual transferability within models, which allows the models to better transfer task-solving capabilities to low-resource languages in the Task-Transfer stage.
MGSM Bn De En Es Fr Ja Ru Sw Te Th Zh LR. HR. Avg.
LinguaLIFT 52.8 62.8 64.8 65.2 55.6 48.4 61.2 53.2 50.4 56.0 54.8 54.0 59.0 57.5
w/o language-align 41.6 56.8 67.2 60.4 58.8 38.4 56.4 45.6 29.2 42.8 46.4 39.8 54.9 49.4
w/o task-transfer 3.20 12.8 14.6 12.4 12.0 9.20 11.2 2.40 1.60 3.60 10.8 2.60 11.9 8.49
Mono-SFT 6.00 58.0 64.8 55.6 55.6 37.6 52.4 4.40 0.00 4.80 40.80 3.80 52.1 34.6
Table 14: Two-stage Training Ablation Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource.
MSVAMP Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
LinguaLIFT 51.3 62.8 67.2 59.9 62.7 54.3 59.7 56.5 55.2 56.0 54.3 60.4 58.6
w/o language-align 36.9 58.7 63.6 58.8 62.1 53.5 55.3 40.8 35.4 53.1 37.7 57.9 51.8
w/o task-transfer 3.8 10.4 16.5 11.7 15.3 7.7 12.0 2.2 4.1 13.7 3.37 12.5 9.74
Mono-SFT 15.2 59.5 64.9 62.9 61.7 52.4 58.3 15.2 18.3 53.7 16.2 59.1 46.2
Table 15: Two-stage Training Ablation Experimental Results on the MSVAMP Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource.
D.2 Ablation of Trainable Modules
To further demonstrate the essential design of the two-stage training approach, we conduct an ablation study on the trainable modules during each stage. This study is performed on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of training each module at the appropriate stage. The detailed results are presented in Tables 16 and 17 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of a specific operation.
The results reveal that performance on high-resource languages suffers when the LLM is frozen during the Language-Align stage. We attribute this decline to catastrophic forgetting, where the initial training on low-resource languages interferes with the model’s ability to retain its high-resource language capabilities. Specifically, the average performance on high-resource languages decreases by 3.5% on the MGSM dataset and 2.6% on the MSVAMP dataset.
When the LLM training is ablated in the Task-Transfer stage, performance significantly deteriorates across all languages. This indicates that relying solely on the language alignment layer is insufficient for learning high-level reasoning tasks and effectively transferring knowledge to low-resource languages. The average performance across all languages drops to 35.4% on the MGSM dataset and 37.8% on the MSVAMP dataset.
Furthermore, freezing the language alignment layer during the Task-Transfer stage results in a decline in performance on low-resource languages. This suggests that the alignment learned in the language alignment layer is disrupted during continuous training, as the average performance on low-resource languages drops by 9.1% on the MGSM dataset and 5.3% on the MSVAMP dataset.
These findings emphasize the benefit of training the language alignment layer first, followed by the LLM training, to improve performance on both low-resource and high-resource languages.
MGSM Bn De En Es Fr Ja Ru Sw Te Th Zh LR. HR. Avg.
LinguaLIFT 52.8 62.8 64.8 65.2 55.6 48.4 61.2 53.2 50.4 56.0 54.8 54.0 59.0 57.5
w/o freezing LLM in Language-Align 50.6 59.3 62.0 60.1 52.7 45.0 58.2 52.6 48.2 54.5 51.5 51.5 55.5 54.1
w/o training LLM in Task-Transfer 26.4 42.4 47.6 45.6 40.4 30.4 41.2 26.8 24.8 28.4 35.6 26.6 40.5 35.4
w/o freezing Language Alignment Layer in Task-Transfer 46.8 60.2 66.4 59.2 57.8 46.8 58.8 45.2 39.6 48.0 49.6 44.9 57.0 52.6
Table 16: Trainable Modules Ablation Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource.
MSVAMP Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
LinguaLIFT 51.3 62.8 67.2 59.9 62.7 54.3 59.7 56.5 55.2 56.0 54.3 60.4 58.6
w/o freezing LLM in Language-Align 50.7 59.9 65.4 57.4 55.6 50.1 57.5 55.6 54.9 53.0 53.7 57.0 56.0
w/o training LLM in Task-Transfer 29.3 49.7 51.3 45.6 47.4 28.8 35.8 27.1 24.8 38.3 27.1 42.4 37.8
w/o freezing Language Alignment Layer in Task-Transfer 42.4 60.4 65.6 60.2 60.8 53.3 58.5 53.1 51.6 50.8 49.0 58.8 55.9
Table 17: Trainable Modules Ablation Experimental Results on the MSVAMP Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource.
D.3 Analysis of Code-Switch Tuning
Part-of-Speech #Nums
Noun 84726
Verb 42451
Adposition 37727
Pronoun 25012
Adjective 21790
Conjunction 20764
Auxiliary 16409
Adverb 9150
Table 18: Part-of-speech statistics results from the MGSM English test set.
In our experiment, we aimed to scrutinize the influence of the part-of-speech (POS) of the substitution words on the multilingual reasoning performance in the context of code-switched tuning. This analysis is essential as it provides insights into the role of various word categories and their combinations in the effectiveness of language model tuning and reasoning capabilities.
Our initial step involved an examination of the distribution of part-of-speech (POS) tags in the reasoning queries, the results of which are tabulated in Table 18. This analysis facilitated a preliminary understanding of the prevalence of different word classes within the reasoning queries, setting the stage for further investigation into their impact on multilingual reasoning.
Subsequently, we categorized the POS combinations into three distinct groups for a more granular analysis. The first group comprised individual POS categories, specifically Verbs, Adpositions, and Pronouns. These categories were selected due to their fundamental role in sentence construction and their potential to affect the meaning and structure of statements significantly.
The second group consisted of syntactic function combinations, including Verb+Adverb, Adjective+Adverb, and Pronoun+Auxiliary+Conjunction combinations. These combinations were chosen based on their syntactic roles and capacity to influence sentence structure and meaning. They are integral to creating complex sentence structures and are often pivotal in conveying nuanced meanings.
The third group focused on key syntactic structures: Subject-Verb (Noun+Verb) and Prepositional Phrases (Adposition+Noun). These structures were selected due to their central role in sentence construction and their potential to encapsulate core semantic information. They form the backbone of many sentence structures and play a crucial role in interpreting a sentence’s meaning.
By analyzing the impact of these POS categories and combinations on the performance of code-switched tuning, we aim to provide a comprehensive understanding of the interplay between syntax and semantics in the context of multilingual reasoning. This analysis will shed light on the importance of different word classes and their combinations in improving the efficacy of multilingual language models.
Further experiments were conducted to investigate the effects of replacing different types of words (e.g., nouns, verbs, prepositions, etc.) on the multilingual reasoning capabilities of LinguaLIFT models. As indicated in Figure 6, we observed that replacing nouns significantly impacted the model’s reasoning performance within the individual POS categories. This finding leads to the hypothesis that reasoning in query sentences may be closely tied to noun-based understanding, possibly due to the central role nouns play in representing key entities and concepts in the task.
Additionally, we found that syntactic structures involving Subject-Verb combinations and Prepositional Phrases performed the best, surpassing all other POS combinations. This suggests that core arguments (subjects and verbs) and their relational elements (prepositions) are crucial for capturing the key relationships in reasoning tasks.
In contrast, the Adjective+Adverb combination showed the weakest performance, indicating that modifiers are less critical for reasoning tasks. This supports the rationale of previous work (Gaur and Saunshi, 2023; Mirzadeh et al., 2024) that proposed transforming mathematical word problems into symbolic reasoning tasks.
Overall, these findings highlight that, for reasoning tasks, word alignment—particularly involving nouns and verbs—is sufficient to enable the model to generalize reasoning abilities across multiple languages.
Appendix E Supplementary Experiments
E.1 Quantitative Analysis of the Correlation between Multilingual Alignment and Reasoning Performance
To evaluate the cross-lingual alignment capabilities of LLMs, we employ the multilingual sentence retrieval benchmark Tatoeba (Artetxe and Schwenk, 2019), a widely used dataset for assessing ML-LMs. The dataset includes up to 1,000 sentences per language, along with their English translations. In our study, we focus on a subset of the original benchmark that aligns with the language categories used in the reasoning test sets MGSM, ensuring consistency with our experimental design.
We adopt the evaluation procedure outlined in XTREME (Hu et al., 2020) to evaluate cross-lingual alignment. For each language pair, we compute the cosine similarity between sentences in the source language and their closest counterparts in the target language. Following the methods of Yang et al. (2021) and Xie et al. (2022), we use top-1 retrieval accuracy as a metric to quantify alignment between the two languages. A higher retrieval accuracy score indicates a stronger alignment.
(a) MGSM
(b) MSVAMP
Figure 7: Performances on MGSM and MSVAMP test set, where languages are sorted by decreasing retrieval accuracy. The trend indicates the mathemtical reasoning performance worsens along with the decreasing cross-lingual alignment degree.
Test set MGSM MSVAMP
Resource-Level Low-Resource High-Resource Low-Resource High-Resource
$\rho$ 0.733* 0.778* 0.763* 0.783*
Table 19: Spearman’s rank correlation coefficient ($\rho$) between retrieval accuracy (%) on Tatoeba and reasoning accuracy on MGSM. An asterisk (*) signifies a statistically significant correlation (p-value < 0.05).
Figure 7 illustrates that the performance of low-resource languages lags significantly behind that of other languages in both mathematical reasoning tasks and the cross-lingual sentence retrieval task. It also shows that a higher alignment degree tends to correlate with improved mathematical reasoning performance. Additionally, we compute the Spearman’s rank correlation between the retrieval accuracy scores and reasoning performance, as presented in Table 19, which reveals a strong correlation between the two. Both the observed trend and the correlation coefficient confirm that the consistency of multilingual mathematical reasoning is closely tied to the cross-lingual alignment degree.
E.2 Language Transferability in Language Families and Writing Systems
Figure 8: The experimental results grouped by diverse language families and writing systems of different models.
In our analysis, we considered the influence of language families and writing systems on LinguaLIFT’s performance. This approach allowed us to delve into the nuances of language transferability, particularly in low-resource languages that share linguistic characteristics with English.
We organized the MMWP benchmark results according to language family and writing system, following the categorization proposed by Zhang et al. (2023). As depicted in Figure 8, and detailed results of specific low-resource languages reasoning tasks presented in Table 20, LinguaLIFT demonstrates superior performance over all baseline models across different language families. Notably, it excels in languages from the Indo-European-Indo-Iranian, Indo-European-Germanic, and Dravidian families. The Indo-European-Germanic languages, which share a common lineage with English, register stronger performance, underscoring the potential benefits of shared ancestry in cross-lingual transfer. In contrast, languages that diverge significantly from English in terms of syntax, such as those from the Dravidian family, present a more challenging scenario, yet LinguaLIFT still yields impressive results.
In Figure 8 and in Table 21, we also observed similar trends in the context of writing system transfer. LinguaLIFT exhibits exceptional performance in languages that employ shared scripts, particularly those using Latin-based orthographies. This observation underscores the critical role of script similarity in facilitating the transfer of knowledge across languages. The shared orthographic features can provide an additional layer of shared structure that aids in the cross-lingual transfer process, enabling more effective learning and translation across related languages.
In summary, our findings underscore the impressive transferability of LinguaLIFT across different low-resource languages from various language families and writing systems. The method significantly enhances cross-lingual transfer, particularly when linguistic factors such as language family and writing system align with English ones. These results highlight the potential of LinguaLIFT as a robust approach for multilingual processing, capable of leveraging shared linguistic characteristics for improved performance, even in low-resource settings.
Language Family Language MetaMath-7B MathOctopus-Parallel-7B MindMerger AlignIFT-MetaMath
Indo-European-Germanic af 32.4 19.1 40.2 46.4
is 14.3 10.5 35.6 41.9
lb 14.6 8.75 37.6 43.4
Average 20.4 12.8 37.8 43.9
Indo-European-Slavic be 17.9 14.6 37.1 39.6
mk 30.1 15.9 40.7 42.9
sk 32.3 16.3 38.6 42.5
Average 26.8 15.6 38.8 41.7
Indo-European-Indo-Iranian bn 9.86 20.1 37.6 39.5
gu 4.81 6.17 35.8 39.0
hi 24.5 11.1 39.4 41.7
mr 9.62 6.78 34.3 32.6
ne 8.38 6.54 37.9 38.6
Average 11.4 10.1 37.0 38.3
Dravidian ta 6.29 3.95 35.0 36.5
te 6.17 10.2 37.1 36.9
kn 6.41 5.43 33.8 33.4
ml 5.18 5.92 37.4 36.7
Average 6.01 6.38 35.8 35.9
Other ar 23.6 8.63 40.2 42.5
eu 6.29 5.55 35.4 35.6
ha 5.92 5.55 29.1 33.4
hy 5.80 6.04 34.7 39.7
sw 7.52 21.7 36.4 47.6
th 13.4 21.3 38.7 41.7
Average 10.4 11.5 35.8 40.1
Table 20: The detailed experimental results of LinguaLIFT on MMWP benchmark grouped by language family.
Writing Scripts Language MetaMath-7B MathOctopus-Parallel-7B MingMerger AlignIFT-MetaMath
Latin af 32.4 19.1 40.2 46.4
eu 6.29 5.55 35.4 35.6
is 14.3 10.5 35.6 41.9
lb 14.6 8.8 37.6 43.4
sw 7.52 21.7 36.4 47.6
sk 32.3 16.3 38.6 42.5
Average 17.9 13.7 37.3 42.9
Arabic ar 23.6 8.63 40.2 42.5
ha 5.92 5.55 29.1 33.4
Average 14.8 7.09 34.7 38.0
Cyrillic be 17.8 14.6 37.1 39.6
mk 30.1 15.9 40.7 42.9
Average 24.0 15.3 38.9 41.3
Devanagari hi 24.5 11.1 39.4 41.7
mr 9.6 6.8 34.3 32.6
ne 8.4 6.5 37.9 38.6
Average 14.2 8.1 37.2 37.6
Other ta 6.29 3.95 35.0 36.5
te 6.17 10.2 37.1 36.9
ml 5.18 5.92 37.4 36.7
bn 9.86 20.1 37.6 39.5
hy 5.8 6.04 34.7 39.7
kn 6.41 5.43 33.8 33.4
gu 4.81 6.17 35.8 39.0
th 13.4 21.3 38.7 41.7
Average 7.6 10.8 36.3 38.3
Table 21: The detailed experimental results of LinguaLIFT on MMWP benchmark grouped by writing systems.
E.3 Adapting to different types and scales of LLMs
LinguaLIFT can be flexibly adapted to different LLMs. To validate this, we conducted experiments on Mistral-7B (Jiang et al., 2023) and the larger Llama-2-13B (Touvron et al., 2023). As shown in Tables 22 and 23, LinguaLIFT outperforms various baselines, with average accuracy improvements of at least 1.6% and 0.9% for low-resource languages based on Llama-2-13B and Mistral-7B, respectively. These results demonstrate LinguaLIFT’s potential for broader applicability across LLMs.
Llama-2-13B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
MetaMath* 11.6 64.8 67.2 65.2 65.2 42.8 63.6 7.60 6.40 49.2 8.53 59.7 44.4
Multi-SFT
MathOctopus-Parallel* 35.2 44.4 53.2 48.0 48.4 43.2 47.6 42.8 46.8 48.8 41.6 47.7 45.8
QAlign-MetaMathQA* 38.4 62.0 69.2 67.2 62.4 52.4 64.4 46.0 49.6 59.2 44.7 62.4 57.1
Leveraging External Tools or Models
Translate-En* 34.8 53.6 70.8 62.4 54.0 44.4 45.6 44.4 54.0 58.0 44.4 55.5 52.2
LangBridge* 39.2 55.2 65.2 60.8 54.8 33.6 58.8 42.0 42.8 42.0 41.3 52.9 49.4
MindMerger-Soft* 55.2 65.2 68.8 69.6 63.6 60.0 68.0 56.4 59.6 60.4 57.1 65.1 62.7
our methods
LinguaLIFT 57.6 64.4 67.8 70.4 64.0 59.8 67.2 57.8 60.8 59.4 58.7 64.7 62.9
Table 22: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (citation: LLaMA2), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024b); Huang et al. (2024).
Mistral-7B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
MetaMath* 38.4 70.4 78.0 71.2 70.8 50.8 67.2 16.8 34.8 57.2 30.0 66.5 55.6
Multi-SFT
MathOctopus-Parallel* 44.0 50.0 58.4 53.2 47.2 48.0 49.6 51.6 48.8 51.6 48.1 51.1 50.2
QAlign-MetaMathQA* 45.6 59.2 65.8 63.6 59.8 49.4 60.2 55.2 51.2 57.2 50.7 59.3 56.7
Leveraging External Tools or Models
Translate-En* 54.6 50.4 69.7 58.6 56.7 57.2 64.9 47.7 58.7 63.1 53.7 60.1 58.2
LangBridge* 50.0 68.4 65.6 65.6 68.8 58.4 68.4 47.2 60.0 65.6 52.4 65.8 61.8
MindMerger-Soft* 57.6 69.2 79.2 71.6 69.6 57.2 68.4 53.2 59.6 68.8 56.8 69.1 65.4
our methods
LinguaLIFT 58.4 67.2 77.6 72.1 70.6 56.4 66.8 58.4 63.2 68.2 60.0 68.4 65.9
Table 23: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024b); Huang et al. (2024).
E.4 Impact of Different Language Alignment Layers
In the main experiment, we utilize two layers of MLP as the language alignment layers to transfer multilingual alignment from the pre-trained multilingual encoder into the LLMs. We also conducted experiments to evaluate the performance of LinguaLIFT when ablating different mapping layers, as shown in Table 24. In contrast to the findings of Yoon et al. (2024), the two layers of MLP used in our main experiment achieved the best performance. In comparison, using a linear layer resulted in lower performance, likely due to its limited capacity to effectively adapt the pre-trained multilingual encoder to LLM, attributed to the smaller number of parameters.
Architecture # Params Bn De En Es Fr Ja Ru Sw Te Th Zh LR. HR. Avg.
Linear 4 M 53.1 61.6 63.9 63.8 56.6 46.0 56.6 49.0 44.4 50.6 52.2 49.3 57.2 54.4
2 layers MLP 10 M 54.4 62.0 64.8 63.6 56.8 50.0 60.4 55.6 54.0 57.6 54.0 55.4 58.8 57.6
3 layers MLP 14 M 53.8 62.8 66.6 56.6 58.8 50.0 59.4 53.8 53.6 58.2 53.4 54.9 58.2 57.0
Table 24: The ablation experiments of the selection of language alignment layer results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource.
E.5 Impact of Multilingual Encoder Size on Language Alignment
Figure 9: MGSM accuracy(%) of LinguaLIFT models using different kinds and types of multilingual encoders.
We experimented with various multilingual encoders regarding types and sizes and evaluated the corresponding reasoning performance on the XNLI dataset. Figure 9 illustrates the XNLI performance across five different sizes of the mT5 encoder: 270M (Small), 470M (Base), 820M (Large), 2.2B (XL), and 6.7B (XXL). Additionally, we evaluate five types of multilingual encoders: mT5, mBERT, XLM, XLM-R, and LaBSE. Our findings indicate that performance improves significantly as the encoder size increases from 270M to 2.2B for the mT5 encoder, with diminishing returns observed as the model scales beyond 2.2B to 6.7B. These results suggest that while enlarging the encoder size leads to better performance, there is a point beyond which further scaling offers limited improvements. In terms of encoder types, we observe that encoders with stronger language alignment capabilities yield better performance. Specifically, more powerful multilingual models like LaBSE, which is fine-tuned with parallel corpus, are more effective at enhancing the capabilities of low-resource languages, demonstrating that greater language alignment leads to better transfer and generalization across diverse languages.
E.6 Incorporating multilingual reasoning data in LinguaLIFT
Figure 10: Effects of tuning LLM with mixed supervised data. Generally, incorporating multilingual supervised data into LinguaLIFT can achieve a higher ceiling for low-resource language tasks performance.
In some settings, we observed that incorporating multilingual reasoning data during the first stage can further enhance task transfer performance in the second stage. This is evidenced by the improved performance compared to the pure English-only reasoning task training. In the mixed supervision setting, we first fine-tune the language alignment layer during the Language-Align stage using multilingual reasoning instruction data GSM8KINSTRUCT (Chen et al., 2023) and code-switching translation instruction data, followed by fine-tuning LLM with English-only instruction data, MetaMathQA. The experimental results on MGSM are presented in Figure 10. We find that incorporating additional multilingual supervision yields an average performance gain of 1.0% on low-resource reasoning tasks and allows the model to achieve comparable performance on high-resource reasoning tasks relative to the vanilla setting.
E.7 LinguaLIFT CoT Examples
In this section, we present three examples of CoT reasoning generated by LinguaLIFT models. These examples demonstrate the ability of LinguaLIFT to effectively understand low-resource languages and generate corresponding mathematical reasoning processes using the chain-of-thought reasoning strategy.
Figure 11 illustrates a zero-shot CoT example from the MGSM Bengali dataset. In this case, the vanilla Mono-SFT model produces incorrect reasoning, while the LinguaLIFT model generates a correct reasoning process, ultimately leading to the correct answer. Similar trends are observed in the Thai and Swahili examples, shown in Figures 12 and 13, where the LinguaLIFT model corrects the erroneous reasoning produced by the Mono-SFT model when presented with the same mathematical problems.
Figure 11: Example of zero-shot CoT generated by MetaMath-7B when prompted with an example from MGSM Bengali. Translation of the input and response is provided in brown text.
Figure 12: Example of zero-shot CoT generated by MetaMath-7B when prompted with an example from MGSM Thai. Translation of the input and response is provided in brown text.
Figure 13: Example of zero-shot CoT generated by MetaMath-7B when prompted with an example from MGSM Swahili. Translation of the input and response is provided in brown text.